A typical atomic nucleus is about 5 × 10−15 m in radius. Use the uncertainty principle to place a lower limit on the energy an electron must have if it is to be part of a nucleus. 


1
Expert's answer
2022-03-22T15:48:55-0400

Answer

For uncertainty

"P=\frac{\hbar}{r}\\=2.11*10^{-20}kg-m/s"

Now energy

"E=\sqrt{p^2c^2+m^2c^4}"


Putting all values

E="3.3*10^{-12}J"